OpenStudy (anonymous):

Two numbers have a sum of 125. Their difference is 29. Find the larger number. Select one: a. 48 b. 55 c. 77 d. 83

OpenStudy (anonymous):

c. 77

OpenStudy (perl):

x+y = 125 x-y = 29

OpenStudy (anonymous):

I would pay you if i could @CO_oLBoY

OpenStudy (perl):

if you add the two equations you get 2x = 154

OpenStudy (perl):

x = 154/2

OpenStudy (anonymous):

u don't need to @MarthaHearst :)